Will you PLEASE HELP. I am clueless on how to work these problems.
A cell divides into two identical copies every 4 mintues. How many cells will exist after 5 hours.
The level of thorium in a sample decreases by a factor of one-half every 2 million years. A meteorite is discovered to have only 8.6% of its original thorium remaining. How old is the meteorite?

Thanks for your help.

Answer by stanbon(75887)

A cell divides into two identical copies every 4 mintues. How many cells will exist after 5 hours.
----------------
A(t) = A(0)(2)^(t/4) where t is # of minutes.
Note 5 hours is 5*60 = 300 minutes
------------------------------------
A(300) = 1*(2^(300/4)) = 2^75 = 3.77789x10^22
===================================================

The level of thorium in a sample decreases by a factor of one-half every 2 million years. A meteorite is discovered to have only 8.6% of its original thorium remaining. How old is the meteorite?
-----
A(t) = A(0)(1/2)^(t/2 million) where t is in million years
0.086A(0) = A(o)(1/2)^(t/2)
0.086 = (1/2)^(t/2)
Take the log of both sides to get:
(t/2)*log(1/2) = log(0.086)
---
t = 2
t = 2* 3.5395...
t = 7.079.. million years
=============================
Cheers,
Stan H.